OpenStudy (anonymous):

log2 6+log2 a=log2 14

OpenStudy (anonymous):

6+a=14..........a=8

OpenStudy (phi):

The reason people use logarithms is because you can add instead of multiply. In other words log(A B) = log(A) + log(B) Here \[\log_2 6 + \log_2 a= \log_2 14\] \[\log_2 6a = \log_2 14\] exponentiate both sides 6 a = 14 a= 14/6 = 7/3 See http://www.wolframalpha.com/input/?i=solve+log2+6%2Blog2+a%3Dlog2+14